write the first 5 terms of the following recursive sequence: a1=5 a_n=a_n-1 +2
Answer by stanbon(75887)

write the first 5 terms of the following recursive sequence: a1=5 a_n=a_n-1 +2
-------
a(1) = 5
a(2) = a(1)+2 = 5+2 = 7
a(3) = 7+2 = 9
a(4) = 9+2 = 11
a(5) = 11+2 = 13
========================
Cheers,
Stan H.
=============